A man holds a string 2 meters long, level with the top of his head. The other end of the string is in the ground 120 centimeters from his feet. How tall is the man?
Found 2 solutions by nyc_function, CharlesG2:
Answer by nyc_function(2741)

2 meters is 200 cm.
Let h = height of man
h^2 + 120^2 = 200^2
h^2 + 14,400 = 40,000
h^2 = 40,000 - 14,000
h^2 = 26,000
sqrt{h^2) = sqrt{26,000)
h = 161.245155 cm
NOTE: sqrt = square root
NOTE: 161.245155 cm means the man is about 5.3 feet tall.
NOTE: 161.245155 cm = 1.61245 meters (in case you need the height of the man in terms of meters).

Answer by CharlesG2(834)

A man holds a string 2 meters long, level with the top of his head. The other end of the string is in the ground 120 centimeters from his feet. How tall is the man?
The way I am reading this since we need to figure out the man's height is as follows:
It is a right triangle, with the hypotenuse (the diagonal, the longest side) being the string which is 2 meters long, or 200 cm long.
There is 100 cm in a meter.
One leg of the right triangle is 120 cm long.
The other leg of the right triangle is the man's height.
Pythagorean Theorem: a^2 + b^2 = c^2
a^2 + (120)^2 = (200)^2
a^2 + 14400 = 40000
a^2 = 25600
a = 160 cm = 1.6 meters is how tall the man is